A rectangular floor measures 80 feet by 39 feet. Find the length of the diagonal of the floor
Answer by nerdybill(7384)

Apply Pythagorean theorem:
Let d = diagonal
then
d^2 = 80^2 + 39^2
d^2 = 6400 + 1521
d^2 = 7921
d = 89 feet